OpenStudy (anonymous):

Multiplication of radicals find the value of (I will post it in equation form below)

OpenStudy (anonymous):

\[\sqrt{2/3} × \sqrt{4/5} \times \sqrt{3/4}\]

OpenStudy (anonymous):

\[\sqrt{2/3} * \sqrt{4/5}* \sqrt{3/4}\] is the same as\[\sqrt{2/3 * 4/5 * 3/4}\] so we can solve this as we would normal fractions to give us \[\sqrt{2/5}\]